OpenStudy (anonymous):

f(x)= square root (3x) whats the rational zero(s)? whats the domain?

Parth (parthkohli):

Domain is the thing you are inputting in f, which is x. Equate f(x) with 0 for the second part.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 0 = \sqrt{3x}}\)

OpenStudy (anonymous):

The zero(s) of a function is the value of x such that f(x) = 0. Thus to find the zero(s) of a function, let f(x) = 0 then solve for x. f(x) = sqrt (3x) = = sqrt (3x) = 0. Square both sides, we obtain 3x = 0 or x = 0 Hence 0 is the only zero of the given function. The domain of a function is the values of x such that the function f(x) is defined. Since we cannot have negative number under the square root so that f(x) is a real number, 3x >= 0 or x >=0 Therefore, the domain of the given function is all real numbers greater or equal to 0.

OpenStudy (anonymous):

Typo on the third line. It should have been f(x) = sqrt (3x) = 0 I apologize :)